# Hear the Scene: Audio-Enhanced Text Spotting

### v2

@@ 1 Introduction
Scene text spotting Tang et al. (2022a); Liu et al. (2020a); Xing et al. (2019); Liu et al. (2023, 2020b); Tang et al. (2022c, b) remains a vibrant area of research due to its extensive applications across various domains such as image-based translation, multimedia content analysis, and accessibility features in visual media. The field Liao et al. (2020); Zhao et al. (2024b); feng2023docpedia; Feng et al. (2023) has evolved from focusing primarily on horizontal or multi-oriented text to addressing more complex arbitrarily-shaped text. This evolution has prompted shifts from traditional annotation formats like horizontal rectangles and quadrilaterals to more sophisticated polygonal annotations, as depicted in Figure 1. Traditionally, scene text spotting involved separate processes for text detection and recognition, where both tasks required annotations for training—text recognition used transcription annotations, and text detection depended on geometric annotations like polygons or bounding boxes. The reliance on detailed geometric annotations for training not only escalates the cost but also increases the manpower required, making the process less efficient and scalable. Recent advancements have introduced methods Li et al. (2024b, a); Zhao et al. (2024c, a); Tang et al. (2024b, 2023, a) that directly identify and classify characters or words within texts, which, while reducing some dependencies, still rely heavily on character-level annotations and thus remain costly. Furthermore, the granularity of these annotations makes the task time-consuming and limits the feasibility of large-scale applications on real-world data. For instance, the annotation of text in videos or cluttered urban scenes can be prohibitively expensive and labor-intensive due to the complexity and volume of data. Given these challenges, there is a growing interest in exploring methods that reduce reliance on expensive annotations. Preliminary efforts, such as the single-point annotation strategy proposed by SPTS Liu et al. (2023), have marked a significant step forward but still leave room for improvement. This paper introduces a novel approach that rethinks scene text spotting from an image captioning perspective, where the text’s location is inferred implicitly through visual attention mechanisms, without direct location supervision. We propose a fully end-to-end architecture, termed EchoSpot, which leverages a query-based interaction between text queries and image embeddings to approximate text locations while ensuring robust text recognition. Our contributions are geared towards simplifying the annotation process, reducing costs, and making text spotting more accessible and scalable. By demonstrating that accurate text spotting can be achieved without explicit location annotations and by integrating novel strategies like circular curriculum learning and audio annotations, this work paves the way for future research and practical applications in the field of optical character recognition and beyond.
@@ 2.2 Weakly-Supervised Text Spotting
The high cost and labor intensity of annotations required by fully-supervised methods Wang et al. (2025); Feng et al. (2024); Shan et al. (2024); Sun et al. (2024) have spurred interest in weakly-supervised approaches. These methods aim to reduce or entirely eliminate the need for detailed geometric annotations. Notable among the emerging solutions are systems that use minimalistic annotations, such as single-point annotations proposed by SPTS Liu et al. (2023), which represent text positions with single points, reducing annotation complexity. Other exploratory works found in preprints like arXiv suggest innovative directions such as using image captions to train models that can identify text regions without explicit location data. These methods leverage advancements in deep learning, such as contrastive learning and transformer architectures, to intuit text locations from high-level descriptions or minimal guidance. Despite these advancements, there remains a discernible gap between the fully-supervised and weakly-supervised approaches in terms of accuracy and reliability, particularly in complex real-world scenarios. Our work builds on these foundations, aiming to bridge this gap by proposing a transcription-only supervised method that simplifies the annotation process to just text transcriptions, eliminating the need for any geometric data. This approach not only aligns with the trend toward reducing annotation dependencies but also enhances the scalability and accessibility of text spotting technologies.

### v3

@@ 1 Introduction
Scene text spotting Tang et al. (2022a); Liu et al. (2020a); Xing et al. (2019); Liu et al. (2023, 2020b); Tang et al. (2022c, b) remains a vibrant area of research due to its extensive applications across various domains such as image-based translation, multimedia content analysis, and accessibility features in visual media. The field Liao et al. (2020); Zhao et al. (2024b); feng2023docpedia; Feng et al. (2023) has evolved from focusing primarily on horizontal or multi-oriented text to addressing more complex, arbitrarily-shaped text. This evolution has prompted shifts from traditional annotation formats like horizontal rectangles and quadrilaterals to more sophisticated polygonal annotations, as depicted in Figure 1. Traditionally, scene text spotting involved separate processes for text detection and recognition, where both tasks required annotations for training. Text recognition used transcription annotations, and text detection depended on geometric annotations like polygons or bounding boxes. The reliance on detailed geometric annotations for training not only escalates the cost but also increases the manpower required, making the process less efficient and scalable. Recent advancements have introduced methods Li et al. (2024b, a); Zhao et al. (2024c, a); Tang et al. (2024b, 2023, a) that directly identify and classify characters or words within texts, which, while reducing some dependencies, still rely heavily on character-level annotations and thus remain costly. Furthermore, the granularity of these annotations makes the task time-consuming and limits the feasibility of large-scale applications on real-world data. For instance, the annotation of text in videos or cluttered urban scenes can be prohibitively expensive and labor-intensive due to the complexity and volume of data. Given these challenges, there is a growing interest in exploring methods that reduce reliance on expensive annotations. Preliminary efforts Wang et al. (2024, 2025b); Sun et al. (2025), such as the single-point annotation strategy proposed by SPTS Liu et al. (2023), have marked a significant step forward but still leave room for improvement. This paper introduces a novel approach that rethinks scene text spotting from an image captioning perspective, where the text’s location is inferred implicitly through visual attention mechanisms, without direct location supervision. We propose a fully end-to-end architecture, termed EchoSpot, which leverages a query-based interaction between text queries and image embeddings to approximate text locations while ensuring robust text recognition. Our contributions are geared towards simplifying the annotation process, reducing costs, and making text spotting more accessible and scalable. By demonstrating that accurate text spotting can be achieved without explicit location annotations and by integrating novel strategies like circular curriculum learning and audio annotations, this work paves the way for future research and practical applications in the field of optical character recognition and beyond.
@@ 2.2 Weakly-Supervised Text Spotting
The high cost and labor intensity of annotations required by fully-supervised methods Wang et al. (2025a); Feng et al. (2024); Shan et al. (2024); Sun et al. (2024) have spurred interest in weakly-supervised approaches. These methods aim to reduce or entirely eliminate the need for detailed geometric annotations. Notable among the emerging solutions are systems that use minimalistic annotations, such as single-point annotations proposed by SPTS Liu et al. (2023), which represent text positions with single points, reducing annotation complexity. Other exploratory works found in preprints like arXiv suggest innovative directions such as using image captions to train models that can identify text regions without explicit location data. These methods leverage advancements in deep learning, such as contrastive learning and transformer architectures, to intuit text locations from high-level descriptions or minimal guidance. Despite these advancements, there remains a discernible gap between the fully-supervised and weakly-supervised approaches in terms of accuracy and reliability, particularly in complex real-world scenarios. Our work builds on these foundations, aiming to bridge this gap by proposing a transcription-only supervised method that simplifies the annotation process to just text transcriptions, eliminating the need for any geometric data. This approach not only aligns with the trend toward reducing annotation dependencies but also enhances the scalability and accessibility of text spotting technologies.